Please help me to solve this equation
3x^2 - 2x = 15x - 10

Found 2 solutions by ilana, MathLover1:
Answer by ilana(307)

For quadratic equations, solve it by getting it in the form ax^2+bx+c=0.
So subtract 15x and add 10 to both sides to get 3x^2-17x+10=0. Now either factor into two binomials or use the quadratic formula to solve for x. I'll factor:
(3x-2)(x-5)=0
I got this by putting (3x-__)(x-__) and trying different factor pairs of 10 for the spaces (10 and 1 or 5 and 2).
So 3x-2=0 and x-5=0
x=2/3,5

Answer by MathLover1(20849)